What is the circumfrance of a circle that is 9 3/4 inches in diameter?
Found 2 solutions by stanbon, jim_thompson5910:
Answer by stanbon(75887)

What is the circumfrance of a circle that is 9 3/4 inches in diameter?
---------------------
C = pi*d
----
C = pi*9.75 inches
--------
C = 30.63
==============
Cheers,
Stan H.
--------------

Answer by jim_thompson5910(35256)

C = pi*d

C = pi*(9 + 3/4)

C = pi*(36/4 + 3/4)

C = pi*(39/4)

C = (39/4)*pi is the exact circumference